The calculator says log 7=0.84509804. Is this a reasonable valu? why or why not?
Answer by 303795(562)

The log of a number is the power of 10 that equals the number
ie log 100 = 2 because 10^2=100
Log 10 = 1 because 10^1 =10
Log 1 = 0 because 10^0=1
So your value of 0.8 seems a reasonable answer between 0 and 1
10^0.84509804 = 7